OpenStudy (anonymous):

Let f(x) = x + 8 and g(x) = x – 4. What is f(x) • g(x)? x2 – 12x + 4 x2 + 4x + 4 x2 + 4x – 32 x2 – 4x – 32

OpenStudy (compassionate):

f(x) = x + 8 and g(x) = x - 4, now what is f(x) x g(x) (x + 8)(x - 4) \[x^2 - 4x + 8x - 24 \rightarrow x^2 + 4x - 32\]